# Dual metrics on the boundary of strictly polyhedral hyperbolic 3-manifolds

### v4

@@ 4 Properness
Suppose that there exists $o\in\phi(D)\cap\partial^{+}_{\infty}d\mathbb{S}^{3}_{s}$. Then $\phi^{-1}(o)$ is a closed $\mathcal{T}$-simplicial set. Define $Y:={\rm ost}(\phi^{-1}(o))$. By construction, $\partial Y\subset X$, hence, the restriction of $\phi$ to $\partial Y$ is an injective path-isometry. Let $\Pi$ be the tangent plane to $\partial_{\infty}^{+}d\mathbb{S}^{3}_{s}$ at $o$ and $P:=\phi(\overline{Y})$, where $\overline{Y}$ is the closure of $Y$ in $D$. Due to future-convexity, we have $P\subset\Pi$. We claim that for any $x,y\in\partial Y$ the segments between $o$ and $\phi(x)$, $\phi(y)$ are disjoint. A proof is similar to the one above. Indeed, suppose the contrary, and suppose that $\phi(x)$ is between $o$ and $\phi(y)$. Once again, ${\rm ost}(x)$ and ${\rm ost}(y)$ are disjoint. We pick $r$ such that for all sufficiently large $i$ the $2r$-neighborhood of $y$ in $(D,d_{i})$ is contained in ${\rm ost}(y)$ and pick neighborhoods $U$ of $\phi(x)$ and $U_{i}$ of $\phi_{i}(x)$ as above. Next, consider a vertex $v\in Y\cap\overline{{\rm ost}}(y)$, which exists since $y\in\partial Y$. We have $\phi(y)=0$. Pick a sequence $y_{i}\in D$ such that $\phi_{i}(y_{i})$ belongs to the segment between $\phi_{i}(v)$ and $\phi_{i}(y)$, and $\phi_{i}(y_{i})$ converge to $\phi(x)$. Then $y_{i}$ converge to $y$. For large enough $i$ the $r$-neighborhoods of $y_{i}$ in $(D,d_{i})$ are contained in ${\rm ost}(y)$. Also for large enough $i$ we have $\phi_{i}(y_{i})\in U_{i}$, thus $\phi_{i}({\rm ost}(y_{i}))$ intersects $FP(\phi_{i}(x))$, which is again a contradiction.
@@
In the second case, we have $P={\rm conv}_{s}(\phi(D))\cap\Pi$ (this is because $\Pi$ cannot contain the $\phi$-image of any triangle not from $\overline{Y}$). In particular, $P$ is a compact convex set. As $\partial P$ is a convex curve in $d\mathbb{S}_{s}^{3}$ around a point at infinity in a degenerate plane $\Pi$, it has length $2\pi$. As the restriction of $\phi$ to $\partial Y$ is a path-isometry, the latter also has length $2\pi$. One can also see that its angles on the side to $Y$ are $\pi$. This is shown in [58, Lemma 6.6]. We quickly sketch the argument. Take a vertex $p$ of $P$ and consider a small enough triangular neighborhood $A$ of $p$ in $P$. Let $v:=\phi^{-1}(p)$. We denote by $A_{i}$ triangular neighborhoods of $\phi_{i}(v)$ in $\phi_{i}(\overline{Y})$ converging to $A$. The sets $\phi_{i}^{-1}(A_{i})$ are triangular neighborhoods of $v$ in $\overline{Y}$ with metrics $d_{i}$. Up to taking a subsequence, they converge to a triangle $B\ni v$ in $\overline{Y}$ with the metric $d$. But as $A$ is metrically a degenerate triangle, $B$ is a geodesic segment on the side of $\overline{Y}$ (having angle $\pi$ at $v$). As all cone angles of $d$ are greater than $2\pi$, $\partial Y$ is a simple closed geodesic of length $2\pi$ in $(D,d)$. This is a contradiction.
@@
We need a short preliminary work. Let $\overline{g}$ be a convex cocompact hyperbolic metric on $N$ and $s\in\partial_{\infty}\mathbb{H}^{3}\backslash\Lambda(\overline{g})$. Consider a small horosphere at $s$ such that it does not intersect $\widetilde{C}(\overline{g})$. Let $\mu(s)\in\partial\widetilde{C}(\overline{g})$ be the closest point from $\widetilde{C}(\overline{g})$ to this horosphere. Clearly, $\mu(s)$ is independent on the choice of a horosphere, and determines a map from $\partial_{\infty}N(\overline{g})$ to $\partial C(\overline{g})$ still denoted by $\mu$. When $\overline{g}$ is Fuchsian of the first kind, we modify $\mu$ so that its target is two copies of $C(\overline{g})$, and $\mu$ remembers from which side of $C(\overline{g})$ the point of $\partial_{\infty}N(\overline{g})$ was located. When $\overline{g}$ is Fuchsian of the second kind, we consider the double of $C(\overline{g})$ along its relative boundary as the target of $\mu$ (keeping track on the side where $s$ was located except when $\mu(s)$ is in the relative boundary of $C(\overline{g})$).
@@
Because $d_{i}^{I}$ converge uniformly to $d^{I}$, the length of a smallest curve in each free homotopy class of closed curves on $\partial M$ is uniformly bounded from above in $\{d_{i}^{I}\}$. Due to the hyperbolic Busemann–Feller Lemma [43], this implies that they are also uniformly bounded from above in $\partial C(\overline{g}_{i})$. Theorem 4.6 implies that the injectivity radii of $\partial C(\overline{g}_{i})$ are uniformly bounded from below. Due to the Theorem 4.5, the length of a smallest curve in each free homotopy class is also uniformly bounded in the hyperbolic representatives of conformal structures in $\partial_{\infty}N(\overline{g}_{i})$. Due to Theorem 2.11 and Theorem 4.6, this means that $\overline{g}_{i}$ stays in a compact subset of $\mathcal{CH}(N)$, and thus, up to taking a subsequence, converges to $\overline{g}\in\mathcal{CH}(N)$. ∎
@@ 5.1 More topological preliminaries
Let $\mathcal{PCH}(N)$ be the space of pointed convex cocompact hyperbolic metrics on $N$, i.e., the space of pairs $(\overline{g},p)$ where $\overline{g}\in\mathcal{CH}(N)$ and $p\in N(\overline{g})$ is a marked point, considered up to isometry sending the marked point to the marked point and isotopic to the identity. Denote by $\mathcal{PCH}_{j}(N)$ the subset of $\mathcal{PCH}(N)$ where $p$ belongs to the geometric end $E_{j}$ corresponding to the $j$-th component $S_{j}$ of $\partial M$. First, we introduce the structure of a fiber bundle on $\mathcal{PCH}_{j}(N)$ with respect to the forgetful map $\sigma_{j}:\mathcal{PCH}_{j}(N)\rightarrow\mathcal{CH}(N)$. For $\overline{g}\in\mathcal{CH}(N)$ denote by $\psi_{j}:E_{j}\rightarrow\mathbb{R}_{+}$ the distance from the marked point to the convex core (so this function depends on $\overline{g}$, but we will suppress this in notation). It is $C^{1}$ ([17, Lemma 1.3.6]) and the level surfaces are strictly convex.
@@
For $\rho^{\prime}\in\widetilde{U}$ we denote by $\widetilde{C}(\rho^{\prime})$ the convex hull of its limit set. Provided that $\widetilde{U}$ is sufficiently small, all surfaces $\widehat{\rm dev}(\rho^{\prime},D_{j})$ are at positive distance from $\widetilde{C}(\rho^{\prime})$. We denote by $\widetilde{E}(\rho^{\prime})$ the connected component of $\mathbb{H}^{3}\backslash\widetilde{C}(\rho^{\prime})$ containing this surface, and by $A$ we denote the set of pairs $(\rho^{\prime},p):\rho^{\prime}\in\widetilde{U},p\in\widetilde{E}(\rho^{\prime})$. Consider the function $\psi_{A}:A\rightarrow\mathbb{R}_{+}$ sending $(\rho^{\prime},p)$ to the distance from $p$ to $\widetilde{C}(\rho^{\prime})$, and the map $A\rightarrow\mathbb{H}^{3}$ sending $(\rho^{\prime},p)$ to the closest point on $\widetilde{C}(\rho^{\prime})$ to $p$. We claim that these maps are continuous. Indeed, when a sequence $\rho_{i}$ converges to $\rho^{\prime}$ in $\widetilde{U}$, the sets $\widetilde{C}(\rho_{i})$ converge to $\widetilde{C}(\rho^{\prime})$ in the Hausdorff sense, see [52, Lemma 3.27]. This easily implies the claim. We now define a map ${\rm dev}:\widetilde{U}\times D_{j}\times\mathbb{R}_{+}\rightarrow\mathbb{H}^{3}$ that sends $(\rho^{\prime},x,r)$ to the point on the same gradient line of $\psi_{A}(\rho^{\prime},.)$ as $\widehat{\rm dev}(\rho^{\prime},x)$ and that is at distance $r$ from $\widetilde{C}(\rho^{\prime})$. By our claim, this map is continuous.
@@
It was shown in [26, Lemma 3.1] that it is a homeomorphism. We fix such a map (which depends only on the choice of $\widehat{V}$).
@@
The group of deck transformations of $\widetilde{S_{\ast V}}\rightarrow S_{\ast V}$ is naturally isomorphic to the pure braid group $B_{0}(\partial M,V)=H_{0}(\partial M,V)/H_{0}^{\sharp}(\partial M,V)$, see [26, Remark 3.5]. Thereby, it is also the group of deck transformations of the covering map $\overline{\mathcal{P}}^{\sharp}(M,V)\rightarrow\overline{\mathcal{P}}(M,V)$. It is also the group of deck tranformations of the covering map $\overline{\mathcal{D}}^{\sharp}_{M}(\partial M,V)\rightarrow\overline{\mathcal{D}}_{M}(\partial M,V)$. By construction, the map $\mathcal{I}^{\sharp}_{V}$ is equivariant with respect to $B_{0}(\partial M,V)$.
@@ 5.2 Weak connectivity properties
Denote by $\mathcal{D}^{\sharp}_{cl}(S,V)$ the space of isotopy classes of concave large spherical cone-metrics $d$ on $(S,V)$ with $V(d)=V$. Lemma 5.5 was first proven in [56, Section 9] for the spaces $\mathcal{D}^{\sharp}_{cl}(S,V)$ when $S$ is the 2-sphere. Then Lemmas 5.5 and 5.6 for $\mathcal{D}^{\sharp}_{cl}(S,V)$ and arbitrary closed $S$ appeared in a manuscript [61] of Schlenker. Brief proof sketches were given there. Basically, Schlenker noticed that a proof of Lemma 5.5 can be given exactly in the same way as it was done for spheres in [56]. For Lemma 5.6 Schlenker remarks that a proof is an easy generalization of the proof of Lemma 5.5. We claim that these results hold straightforwardly also for the spaces $\mathcal{D}^{\sharp}_{M}(\partial M,V)$. Indeed, we only need to modify the proofs to gain control of the smallest geodesic that is contractible not just in $\partial M$, but in $M$. For a metric $d$ on $S$ by ${\rm ml}(d)$ we denote the infimum of lengths of closed geodesics in $(S,d)$ that are contractible in $S$. The control on ${\rm ml}(d)$ in [56] follows from two facts. The first is that if $\lambda>0$ is a number and $d$ is a Riemannian metric on $S$, then ${\rm ml}(\lambda d)=\lambda{\rm ml}(d)$, which clearly holds also for the function ${\rm ml}_{M}$. The second fact is that ${\rm ml}$ is upper-semicontinuous in the Lipschitz topology. For ${\rm ml}_{M}$ this was shown in Corollary 2.8.
@@
It remains to make few easy remarks on the topological boundary of $\mathcal{D}_{c}(\partial M,V)$ in $\mathcal{D}(\partial M,V)$. Let $d^{\sharp}\in\mathcal{D}^{\sharp}(\partial M,\mathcal{T})$ and $\lambda\in\mathbb{R}$ be a number. We denote by $(\lambda*_{\mathcal{T}}d^{\sharp})$ a metric obtained by changing the length $l_{e}$ of each edge $e\in E(\mathcal{T})$ to $e^{\lambda}l_{e}$ if the lengths remain staying $<\pi$ under this operation (otherwise the operation is not defined). We will use some corollaries of the following easy lemma
@@
Let $T$ be a spherical triangle with side lengths $a$, $b$ and $c$ less than $\pi$. Consider a spherical triangle $T^{\prime}$ with the lengths $e^{\lambda}a$, $e^{\lambda}b$, $e^{\lambda}c$ for $\lambda>0$ such that the lengths of the new triangle are $<\pi$. Then the angles of $T^{\prime}$ are bigger than the respective angles of $T$.
The proof is identical to the proof of a similar lemma for hyperbolic triangles ([50, Lemma 2.3.9]) with all inequalities reversed.

### v5

@@ 4 Properness
Suppose that there exists $o\in\phi(D)\cap\partial^{+}_{\infty}d\mathbb{S}^{3}_{s}$. Then $\phi^{-1}(o)$ is a closed $\mathcal{T}$-simplicial set. Define $Y:={\rm ost}(\phi^{-1}(o))$. By construction, $\partial Y\subset X$, hence, the restriction of $\phi$ to $\partial Y$ is an injective path-isometry. Let $\Pi$ be the tangent plane to $\partial_{\infty}^{+}d\mathbb{S}^{3}_{s}$ at $o$ and $P:=\phi(\overline{Y})$, where $\overline{Y}$ is the closure of $Y$ in $D$. Due to future-convexity, we have $P\subset\Pi$. We claim that for any $x,y\in\partial Y$ the segments between $o$ and $\phi(x)$, $\phi(y)$ are disjoint. A proof is similar to the one above. Indeed, suppose the contrary, and suppose that $\phi(x)$ is between $o$ and $\phi(y)$. Once again, ${\rm ost}(x)$ and ${\rm ost}(y)$ are disjoint. We pick $r$ such that for all sufficiently large $i$ the $2r$-neighborhood of $y$ in $(D,d_{i})$ is contained in ${\rm ost}(y)$ and pick neighborhoods $U$ of $\phi(x)$ and $U_{i}$ of $\phi_{i}(x)$ as above. Next, consider a vertex $v\in Y\cap\overline{{\rm ost}}(y)$, which exists since $y\in\partial Y$. We have $\phi(v)=o$. Pick a sequence $y_{i}\in D$ such that $\phi_{i}(y_{i})$ belongs to the segment between $\phi_{i}(v)$ and $\phi_{i}(y)$, and $\phi_{i}(y_{i})$ converge to $\phi(x)$. Then $y_{i}$ converge to $y$. For large enough $i$ the $r$-neighborhoods of $y_{i}$ in $(D,d_{i})$ are contained in ${\rm ost}(y)$. Next, for $y^{\prime}$ in the $r$-neighborhood of $y_{i}$, the space-like distance between $\phi_{i}(y^{\prime})$ and $\phi_{i}(y_{i})$ is at least the intrinsic distance between $y^{\prime}$ and $y_{i}$. Also for large enough $i$ we have $\phi_{i}(y_{i})\in U_{i}$, thus $\phi_{i}({\rm ost}(y_{i}))$ intersects $FP(\phi_{i}(x))$. Because ${\rm ost}(y_{i})\subset{\rm ost}(y)$, $\phi_{i}({\rm ost}(y))$ intersects $FP(\phi_{i}(x))$, which is again a contradiction.
@@
In the second case, we have $P={\rm conv}_{s}(\phi(D))\cap\Pi$ (this is because $\Pi$ cannot contain the $\phi$-image of any triangle not from $\overline{Y}$). In particular, $P$ is a compact convex set. As $\partial P$ is a convex curve in $d\mathbb{S}_{s}^{3}$ around a point at infinity in a degenerate plane $\Pi$, it has length $2\pi$. As the restriction of $\phi$ to $\partial Y$ is a path-isometry, the latter also has length $2\pi$. One can also see that its angles on the side to $Y$ are $\pi$. This is shown in [61, Lemma 6.6]. We quickly sketch the argument. Take a vertex $p$ of $P$ and consider a small enough triangular neighborhood $A$ of $p$ in $P$. Let $v:=\phi^{-1}(p)$. We denote by $A_{i}$ triangular neighborhoods of $\phi_{i}(v)$ in $\phi_{i}(\overline{Y})$ converging to $A$. The sets $\phi_{i}^{-1}(A_{i})$ are triangular neighborhoods of $v$ in $\overline{Y}$ with metrics $d_{i}$. Up to taking a subsequence, they converge to a triangle $B\ni v$ in $\overline{Y}$ with the metric $d$. But as $A$ is metrically a degenerate triangle, $B$ is a geodesic segment on the side of $\overline{Y}$ (having angle $\pi$ at $v$). As all cone angles of $d$ are greater than $2\pi$, $\partial Y$ is a simple closed geodesic of length $2\pi$ in $(D,d)$. This is a contradiction.
@@
We need a short preliminary work. Let $\overline{g}$ be a convex cocompact hyperbolic metric on $N$ and $s\in\partial_{\infty}\mathbb{H}^{3}\backslash\Lambda(\overline{g})$. Consider a small horosphere at $s$ such that it does not intersect $\widetilde{C}(\overline{g})$. Let $\mu(s)\in\partial\widetilde{C}(\overline{g})$ be the closest point from $\widetilde{C}(\overline{g})$ to this horosphere. Clearly, $\mu(s)$ is independent on the choice of a horosphere, and determines a map from $\partial_{\infty}N(\overline{g})$ to $\partial C(\overline{g})$, still denoted by $\mu$. When $\overline{g}$ is Fuchsian of the first kind, we modify $\mu$ so that its target is two copies of $C(\overline{g})$, and $\mu$ remembers, from which side of $C(\overline{g})$ the point of $\partial_{\infty}N(\overline{g})$ was located. When $\overline{g}$ is Fuchsian of the second kind, we consider the double of $C(\overline{g})$ along its relative boundary as the target of $\mu$ (keeping track on the side where $s$ was located except when $\mu(s)$ is in the relative boundary of $C(\overline{g})$).
@@
Because $d_{i}^{I}$ converge uniformly to $d^{I}$, the length of a smallest curve in each free homotopy class of closed curves on $\partial M$ is uniformly bounded from above in $\{d_{i}^{I}\}$. Due to the hyperbolic Busemann–Feller Lemma [45], this implies that they are also uniformly bounded from above in $\partial C(\overline{g}_{i})$. Theorem 4.7 implies that the injectivity radii of $\partial C(\overline{g}_{i})$ are uniformly bounded from below. Due to Theorem 4.6, the length of a smallest curve in each free homotopy class is also uniformly bounded in the hyperbolic representatives of conformal structures in $\partial_{\infty}N(\overline{g}_{i})$. Due to Theorem 2.12 and Theorem 4.7, this means that $\overline{g}_{i}$ stay in a compact subset of $\mathcal{CH}(N)$, and thus, up to taking a subsequence, converge to $\overline{g}\in\mathcal{CH}(N)$. ∎
@@ 5.1 More topological preliminaries
Let $\mathcal{PCH}(N)$ be the space of pointed convex cocompact hyperbolic metrics on $N$, i.e., the space of pairs $(\overline{g},p)$ where $\overline{g}\in\mathcal{CH}(N)$ and $p\in N(\overline{g})$ is a marked point, considered up to isometry sending the marked point to the marked point and isotopic to the identity. Denote by $\mathcal{PCH}_{j}(N)$ the subset of $\mathcal{PCH}(N)$ where $p$ belongs to the geometric end $E_{j}$ corresponding to the $j$-th component $S_{j}$ of $\partial M$. First, we introduce the structure of a fiber bundle on $\mathcal{PCH}_{j}(N)$ with respect to the forgetful map $\sigma_{j}:\mathcal{PCH}_{j}(N)\rightarrow\mathcal{CH}(N)$. For $\overline{g}\in\mathcal{CH}(N)$ denote by $\psi_{j}:E_{j}\rightarrow\mathbb{R}_{+}$ the distance from the marked point to the convex core (so this function depends on $\overline{g}$, but we will suppress this in notation). It is $C^{1}$ ([19, Lemma 1.3.6]) and the level surfaces are strictly convex.
@@
For $\rho^{\prime}\in\widetilde{U}$ we denote by $\widetilde{C}(\rho^{\prime})$ the convex hull of its limit set. Provided that $\widetilde{U}$ is sufficiently small, all surfaces $\widehat{\rm dev}(\rho^{\prime},D_{j})$ are at positive distance from $\widetilde{C}(\rho^{\prime})$. We denote by $\widetilde{E}(\rho^{\prime})$ the connected component of $\mathbb{H}^{3}\backslash\widetilde{C}(\rho^{\prime})$ containing this surface, and by $A$ we denote the set of pairs $(\rho^{\prime},p):\rho^{\prime}\in\widetilde{U},p\in\widetilde{E}(\rho^{\prime})$. Consider the function $\psi_{A}:A\rightarrow\mathbb{R}_{+}$ sending $(\rho^{\prime},p)$ to the distance from $p$ to $\widetilde{C}(\rho^{\prime})$, and the map $A\rightarrow\mathbb{H}^{3}$ sending $(\rho^{\prime},p)$ to the closest point on $\widetilde{C}(\rho^{\prime})$ to $p$. We claim that these maps are continuous. Indeed, when a sequence $\rho_{i}$ converges to $\rho^{\prime}$ in $\widetilde{U}$, the sets $\widetilde{C}(\rho_{i})$ converge to $\widetilde{C}(\rho^{\prime})$ in the Hausdorff sense, see [54, Lemma 3.27]. This easily implies the claim. We now define a map ${\rm dev}:\widetilde{U}\times D_{j}\times\mathbb{R}_{+}\rightarrow\mathbb{H}^{3}$ that sends $(\rho^{\prime},x,r)$ to the point on the same gradient line of $\psi_{A}(\rho^{\prime},.)$ as $\widehat{\rm dev}(\rho^{\prime},x)$ and that is at distance $r$ from $\widetilde{C}(\rho^{\prime})$. By our claim, this map is continuous.
@@
It was shown in [28, Lemma 3.1] that it is a homeomorphism. We fix such a map (which depends only on the choice of $\widehat{V}$).
@@
The group of deck transformations of $\widetilde{S_{\ast V}}\rightarrow S_{\ast V}$ is naturally isomorphic to the pure braid group $B_{0}(\partial M,V)=H_{0}(\partial M,V)/H_{0}^{\sharp}(\partial M,V)$, see [28, Remark 3.5]. Thereby, it is also the group of deck transformations of the covering map $\overline{\mathcal{P}}^{\sharp}(M,V)\rightarrow\overline{\mathcal{P}}(M,V)$. It is also the group of deck tranformations of the covering map $\overline{\mathcal{D}}^{\sharp}_{M}(\partial M,V)\rightarrow\overline{\mathcal{D}}_{M}(\partial M,V)$. By construction, the map $\mathcal{I}^{\sharp}_{V}$ is equivariant with respect to $B_{0}(\partial M,V)$.
@@ 5.2 Weak connectivity properties
Denote by $\mathcal{D}^{\sharp}_{cl}(S,V)$ the space of isotopy classes of concave large spherical cone-metrics $d$ on $(S,V)$ with $V(d)=V$. Lemma 5.5 was first proven in [58, Section 9] for the spaces $\mathcal{D}^{\sharp}_{cl}(S,V)$ when $S$ is the 2-sphere. Then Lemmas 5.5 and 5.6 for $\mathcal{D}^{\sharp}_{cl}(S,V)$ and arbitrary closed $S$ appeared in a manuscript [64] of Schlenker. Brief proof sketches were given there. Basically, Schlenker noticed that a proof of Lemma 5.5 can be given exactly in the same way as it was done for spheres in [58]. For Lemma 5.6 Schlenker remarks that a proof is an easy generalization of the proof of Lemma 5.5. We claim that these results hold straightforwardly also for the spaces $\mathcal{D}^{\sharp}_{M}(\partial M,V)$. Indeed, we only need to modify the proofs to gain control of the smallest geodesic that is contractible not just in $\partial M$, but in $M$. For a metric $d$ on $S$ by ${\rm ml}(d)$ we denote the infimum of lengths of closed geodesics in $(S,d)$ that are contractible in $S$. The control on ${\rm ml}(d)$ in [58] follows from two facts. The first is that if $\lambda>0$ is a number and $d$ is a Riemannian metric on $S$, then ${\rm ml}(\lambda d)=\lambda{\rm ml}(d)$, which clearly holds also for the function ${\rm ml}_{M}$. The second fact is that ${\rm ml}$ is upper-semicontinuous in the Lipschitz topology. For ${\rm ml}_{M}$ this was shown in Corollary 2.8.
@@
It remains to make a few easy remarks on the topological boundary of $\mathcal{D}_{c}(\partial M,V)$ in $\mathcal{D}(\partial M,V)$. Let $d^{\sharp}\in\mathcal{D}^{\sharp}(\partial M,\mathcal{T})$ and $\lambda\in\mathbb{R}$ be a number. We denote by $(\lambda*_{\mathcal{T}}d^{\sharp})$ a metric obtained by changing the length $l_{e}$ of each edge $e\in E(\mathcal{T})$ to $e^{\lambda}l_{e}$ if the lengths remain staying $<\pi$ under this operation (otherwise the operation is not defined). We will use some corollaries of the following easy lemma
@@
Let $T$ be a spherical triangle with side lengths $a$, $b$ and $c$ less than $\pi$. Consider a spherical triangle $T^{\prime}$ with the lengths $e^{\lambda}a$, $e^{\lambda}b$, $e^{\lambda}c$ for $\lambda>0$ such that the side lengths of the new triangle are $<\pi$. Then the angles of $T^{\prime}$ are bigger than the respective angles of $T$.
The proof is identical to the proof of a similar lemma for hyperbolic triangles ([52, Lemma 2.3.9]) with all inequalities reversed.